# UEMM-Air: A Synthetic Multi-modal Dataset for Unmanned Aerial Vehicle Object Detection

### v1

@@ Abstract
The development of multi-modal object detection for Unmanned Aerial Vehicles (UAVs) typically relies on a large amount of pixel-aligned multi-modal image data. However, existing datasets face challenges such as limited modalities, high construction costs, and imprecise annotations. To this end, we propose a synthetic multi-modal UAV-based object detection dataset, UEMM-Air. Specially, we simulate various UAV flight scenarios and object types using the Unreal Engine (UE). Then we design the UAV’s flight logic to automatically collect data from different scenarios, perspectives, and altitudes. Finally, we propose a novel heuristic automatic annotation algorithm to generate accurate object detection labels. In total, our UEMM-Air consists of 20k pairs of images with 5 modalities and precise annotations. Moreover, we conduct numerous experiments and establish new benchmark results on our dataset. We found that models pre-trained on UEMM-Air exhibit better performance on downstream tasks compared to other similar datasets. The dataset is publicly available (https://github.com/1e12Leon/UEMM-Air) to support the research of multi-modal UAV object detection models.
1 Introduction
@@
With the advancement of Unmanned Aerial Vehicles (UAV) technology [23, 31]and deep learning [15, 30], vision tasks of UAV have shown great potential in many fields such as agriculture and rescue [31, 33]. The object detection from the perspective of UAV (UAV-OD) [22, 37, 42] has attracted widespread attention from research community. This is due to its ability to efficiently complete tasks such as urban monitoring [39] and military reconnaissance [18]. Unlike general object detection [43], UAV-OD tasks exhibit characteristics such as complex backgrounds and small objects. Therefore, models trained on general object detection datasets [17, 8] can hardly be applied to UAV-OD tasks.
@@
To this end, many scholars have constructed object detection datasets from the perspective of UAV. For example, VisDrone [5] and UAVDT [4] consider various scenes, weather conditions, and environments, providing a good benchmark for UAV-OD tasks. However, with the development of multi-modal learning [1, 38], these datasets are facing challenges such as being single-modality and having insufficient data. DroneVehicle [35] utilizes two image modalities: infrared and visible, with the infrared modality enhancing detection accuracy in nighttime scenes. However, due to the request for manual labeling and aligning two modalities images, the dataset annotation cost is relatively high. AU-AIR dataset [3] takes into account the potential value of UAV parameters for object detection. However, it covers relatively only a few scenes and images have some imprecise annotations. These drawbacks are not conducive to model training.
@@
As illustrated in Fig.1, to address the aforementioned issues, we construct a new synthetic multi-modal UAV object detection dataset UEMM-Air (Unreal Engine Multi-modal Dataset for UAV-based Object Detection). Specifically, we first utilize the Unreal Engine (UE) [7] and AirSim [29] framework to build various simulated scenarios for UAV flights. Subsequently, we implement automatic UAV flight control and collect data at different altitudes, scenes, and modalities. Finally, we design an annotation algorithm to automatically generate object detection labels. It is worth noting that our dataset contains five modalities, including visible, depth, segmentation, surface normals, and UAV IMU parameters. We also provide more accurate annotation information, including fine-grained and scene labels, which are beneficial for various types of object detection tasks. Additionally, we conduct experiments on multiple types of tasks in the field of object detection, demonstrating the research significance of our dataset across various tasks. The main contributions are as follows:
We propose a synthetic multi-modal UAV-based object detection dataset, UEMM-Air. It includes a greater number of modalities, more accurate annotations, richer scenes, and fine-grained labels than existing datasets (e.g. SynDrone). To the best of our knowledge, UEMM-Air is the UAV-OD dataset with the largest number of modalities.
@@
We introduce a new heuristic algorithm for automatic data annotation. Compared with labeling strategies in SynDrone, ours can provide more accurate annotations by introducing segmentation and depth modalities to enhance the identification of objects. Especially when it comes to addressing visually overlapped objects.
We conduct experiments on multiple types of tasks in the field of object detection, providing new benchmark results. Due to our accurate labeling and rich modalities, our UEMM-Air provides remarkable transferability for downstream tasks.
@@
| Dataset | Year | MM | # modalities | # images | # classes | Size [px] | Angle |
| Stanford-Drone [28] | 2016 | ✗ | - | - | 7 | 1450×1080 | 90 |
| UAVDT [4] | 2018 | ✗ | - | 40376 | 3 | 1080×540 | variable |
| VisDrone [5] | 2018 | ✗ | - | 8629 | 10 | 1920×1080 | variable |
| AU-AIR [3] | 2020 | ✓ | 2 | 32823 | 8 | 1920×1080 | 45 to 90 |
| Drone-Vehicle [34] | 2022 | ✓ | 2 | 28 | 5 | 640×512 | 90 |
| SynDrone [27] | 2023 | ✓ | 3 | 60 | 9 | 1920×1080 | 30,60,90 |
| UEMM-Air (Ours) | 2024 | ✓ | 5 | 20 | 13 | 1920×1080 | variable |
2 Related Work
2.1 UAV-based Object Detection
@@
Due to the typically top-down perspective of UAVs and the relatively small scale of the objects, general object detection methods [9, 25, 24, 2, 20] tend to be not suitable for UAV-based Object Detection (UAV-OD) tasks. The mainstream methods for UAV-based object detection primarily employ coarse-to-fine strategies [6, 16, 40]. Initially, the detection process focuses on identifying larger objects, while concurrently pinpointing dense subregions containing small objects. These subregions are subsequently utilized as inputs for the model to refine detection results. For example, a CZDetector [21]employed a density-based cropping algorithm to identify regions with crowded objects and then increased the size of those regions to enhance the training dataset. Alternatively, Koyun et al. [14] utilizes a Gaussian mixture model to supervise the detector in generating object clusters composed of focusing regions. And DMNet [16] optimizes region selection by conducting a density map guided connected crop generation.
2.2 UAV-OD Datasets
@@
The UAV-based object detection dataset provides multi-class images and videos captured by UAVs. These datasets are significant for promoting the research and development of various computer vision tasks, including object tracking, path planning, and scene understanding. We summarized several commonly used UAV-OD datasets in Table 1. Due to the high cost of obtaining images taken by UAV, early datasets, such as UAVDT and VisDrone, have only a single modality, namely RGB images. In addition, owing to hardware limitations, datasets like Drone-Vehicle are small, low-resolution, and difficult to expand.
@@
Stanford-Drone [28] is a large-scale dataset containing overhead images and videos of multi-class of objects moving and interacting at Stanford University. This dataset can be used for learning and evaluating multi-object tracking, activity understanding, and trajectory prediction.
@@
AU-AIR [3] includes extracted frames meta-data, bounding box annotations for traffic-related object categories, and multi-modal flight sensor data. The dataset is captured at low altitudes at the intersection and covers various lighting conditions.
@@
3 UEMM-Air Dataset
3.1 Scene Construction and Flight Control Logic
Previous UAV-OD datasets have been limited in scene diversity, which tends to affect model generalization. Therefore, we aim to construct a dataset with richer scenes to improve the performance of the models. To be specific, we utilize Unreal Engine with CityBLD [32] plugin. It can create cities of almost any size and style in a very short time to simulate scenarios in the real world. we build several scenes in Unreal Engine, including cities, parks, highways, etc. We collect a total of 13 categories and more than hundreds of vehicle models.
@@ 3.1 Scene Construction and Flight Control Logic
To collect data, we controll the UAV to fly and take pictures in Unreal Engine. Specifically, we build an Unmanned Aerial Vehicle simulator using AirSim and Pygame. When flying to a satisfactory shooting point, we control the UAV to fly within a height range of 5 meters to 50 meters above the horizontal surface, taking a set of photos every 5 meters up. The camera rotates from 0 degrees to 90 degrees by 5 degrees for each step. To obtain aligned pictures of different modalities, our simulator temporarily stops running when taking photos.
3.2 Automatic Image Segmentation and Annotation
Most of the existing object detection datasets are manually annotated. Manual image annotation faces challenges in terms of accuracy and efficiency, especially when dealing with a large number of labels or low-resolution images.
To avoid manual annotation, the SynDrone dataset employs an automatic image labeling algorithm. They derive the absolute coordinates of UAV and vehicles from Unreal Engine, then obtain the detection frame of the object by analyzing their relative position. However, this strategy causes some incorrect annotations where objects are visibly blocked but their coordinates are still marked on the image, as illustrated in Fig.2.
To avoid mislabeling visually overlapped objects, we utilize depth images where pixel value represents the distance from the object to the camera plane to perform a secondary annotation. Intuitively, depth values mildly change on each object and a depth value jump indicates multiple objects existences. Therefore overlapped objects can be correctly identified through depth observation. We detect depth mutations within segmented bounding boxes to confirm object edges, adjusting labels accordingly.
3.3 Acquisition Setup
@@
| Method | Fine-tuned | Pre-trained | ||
| YOLOv8 | VisDrone | SynDrone | 25.6% | 16.0% |
| UEMM-Air | 28.3% | 17.7% | ||
| UAV-DT | SynDrone | 85.4% | 55.6% | |
| UEMM-Air | 86.2% | 56.1% | ||
| FaterRCNN | VisDrone | SynDrone | 5.1% | 1.1% |
| UEMM-Air | 5.5% | 2.3% | ||
| UAV-DT | SynDrone | 48.0% | 9.5% | |
| UEMM-Air | 53.8% | 14.7% |
4 Benchmark and Experiments
4.1 Experimental Setup

### v2

# UEMM-Air: Make Unmanned Aerial Vehicles Perform More Multi-modal Tasks

@@ Abstract
The development of multi-modal learning for Unmanned Aerial Vehicles (UAVs) typically relies on a large amount of pixel-aligned multi-modal image data. However, existing datasets face challenges such as limited modalities, high construction costs, and imprecise annotations. To this end, we propose a synthetic multi-modal UAV-based multi-task dataset, UEMM-Air. Specifically, we simulate various UAV flight scenarios and object types using the Unreal Engine (UE). Then we design the UAV’s flight logic to automatically collect data from different scenarios, perspectives, and altitudes. Furthermore, we propose a novel heuristic automatic annotation algorithm to generate accurate object detection labels. Finally, we utilize labels to generate text descriptions of images to make our UEMM-Air support more cross-modality tasks. In total, our UEMM-Air consists of 120k pairs of images with 6 modalities and precise annotations. Moreover, we conduct numerous experiments and establish new benchmark results on our dataset. We also found that models pre-trained on UEMM-Air exhibit better performance on downstream tasks compared to other similar datasets. The dataset is publicly available (https://github.com/1e12Leon/UEMM-Air) to support the research of multi-modal tasks on UAVs.
Index Terms:
Unmanned Aerial Vehicles, Large Scale Dataset, Multi-modal, Multi-taskI Introduction
@@
With the advancement of Unmanned Aerial Vehicles (UAV) technology [1, 2, 3, 4, 5] and deep learning [6, 7], vision perception tasks of UAV have shown great potential in many fields such as urban monitoring, military reconnaissance and rescue [8, 5, 9, 10]. Unlike general vision task [11], UAV tasks exhibit characteristics such as complex backgrounds, varying scales, and small objects. Therefore, models trained on general datasets [12, 13] can hardly be directly applied to UAV platforms. To this end, many scholars have constructed vision datasets from the perspective of UAV. For example, VisDrone [14] and UAVDT [15] consider various scenes, weather conditions, and environments, providing good benchmarks for UAV-based Object Detection (UAV-OD) [16, 17, 18, 19] tasks. SkyScenes [20] encompassing different layouts, weather, and daytime conditions with corresponding dense annotations and viewpoint metadata for UAV-based Segmentation (UAV-Seg) [21, 22]. The aforementioned datasets make significant contributions to traditional UAV vision tasks.
@@
However, with the development of multi-modal learning [23, 24], the above datasets are facing challenges such as being single-modality and having insufficient data. Therefore, the research communities are gradually moving from single-modal to multi-modal tasks. For instance, DroneVehicle [25] utilizes two image modalities: infrared and visible, with the infrared modality enhancing detection accuracy in nighttime scenes. However, due to the request for manual labeling and aligning two modalities images, the dataset annotation cost is relatively high. Although AU-AIR dataset [26] takes into account the potential value of UAV parameters, it covers relatively only a few scenes and its labeled samples have some imprecise annotations. Drawbacks in above datasets are adverse to model training. Furthermore, all the existing datasets are unable to support alignment of vision and language to perform zero-shot inference like the CLIP [27] model in the field of UAVs. Similar to satellite remote sensing, UAVs also need numerous vision-language applications such as open-vocabulary object detection [28], referring image segmentation [29], and multi-modal large language models (MLLMs) [30], etc.
Motivated by the above observations, we construct a new large-scale synthetic UAV environmental perception dataset, UEMM-Air, to facilitate further research on the UAV field. As represented in Fig. 1 and Fig. 2, we assign UEMM-Air three significant characteristics. (1) Multi-modal: our dataset contains six modalities, including visible, depth, segmentation, surface normals, UAV IMU parameters, and captions. (2) Multi-task: our dataset is capable of supporting tasks such as object detection, instance segmentation, image-text contrastive learning, and referring image segmentation, etc. (3) Multi-semantic: our dataset covers a variety of scenarios and perspective, and has fine-grained category information.
@@
To achieve these objectives, we first utilize the Unreal Engine (UE) [31] and AirSim [32] framework to build various simulated scenarios for UAV flights. Subsequently, we implement automatic UAV flight control and collect data at different altitudes, scenes, and modalities. Furthermore, we design an annotation algorithm to automatically generate object detection labels. Finally, we generate text descriptions for different cross-modal tasks according to existing detection and segmentation annotations. We also implement representative and impressive methods to systematically investigate the potential and challenges brought by UEMM-Air. The experimental results demonstrate the research significance of our dataset across various tasks.
The main contributions of this article are as follows:
We propose a new synthetic UAV-based environmental perception dataset, UEMM-Air. To the best of our knowledge, it is the largest dataset in terms of existing data scale, featuring the most paired modality types and the highest variety of task types.
@@
We introduce a new heuristic algorithm for automatic data annotation. Compared with labeling strategies in other synthetic datasets, ours can provide more accurate annotations by introducing segmentation and depth modalities to enhance the identification of objects, especially in addressing visually overlapped objects.
We conduct experiments on multiple types of tasks in the field of UAV, providing new benchmark results. Our UEMM-Air expands the field of UAV environmental perception from purely visual to multi-modal tasks.
The remainder of this paper is structured as follows. In Section II, we review related literature on UAV environmental perception and existing datasets. Section III introduces our UEMM-Air. Specifically, Section III-A introduces the automatic data collection methods we used. Then, in Section III-B, we discuss the configurations and advantages of each modality. We also outline the strategies for automatic annotation generation and the methods for cross-modal generation, respectively in Sections III-C and Section III-D. Next, Section IV validates the accuracy of our annotations and demonstrates the benchmarks of our UEMM-Air across multiple tasks, including object detection, instance segmentation, referring segmentation, and cross-modal retrieval. Finally, Section V presents our discussion of the advantages and limitations of our UEMM-Air and concludes this paper.
@@
| Tasks | Dataset | Year | MM | # modalities | # images | # classes | Size [px] | Angle |
| Det | Stanford-Drone [33] | 2016 | ✗ | - | - | 7 | 1450×1080 | 90 |
| UAVDT [15] | 2018 | ✗ | - | 40376 | 3 | 1080×540 | variable | |
| VisDrone [14] | 2018 | ✗ | - | 8629 | 10 | 1920×1080 | variable | |
| AU-AIR [26] | 2020 | ✓ | 2 | 32823 | 8 | 1920×1080 | 45 to 90 | |
| Drone-Vehicle [34] | 2022 | ✓ | 2 | 28 | 5 | 640×512 | 90 | |
| HIT-UAV [35] | 2023 | ✓ | 2 | 2898 | 4 | 640×512 | 30 to 90 | |
| RTDOD [36] | 2023 | ✓ | 2 | 16192 | 10 | 1280×720 | variable | |
| State-Air [37] | 2024 | ✓ | 2 | 2864 | 4 | 1280×720 | variable | |
| Seg | UAVid [38] | 2020 | ✗ | - | 270 | 8 | 3840×2160 | 45 |
| NWPU_YRCC [39] | 2020 | ✗ | - | 814 | 3 | 1600×640 | variable | |
| RescueNet [40] | 2022 | ✗ | - | 4494 | 10 | 3000×4000 | variable | |
| AFID [41] | 2023 | ✗ | - | 816 | 8 | 2560×1440 | variable | |
| SkyScenes [20] | 2024 | ✓ | 2 | 33.6 | 28 | 2160×1440 | variable | |
| Dets, Seg | SynDrone [42] | 2023 | ✓ | 3 | 60 | 9 | 1920×1080 | 30,60,90 |
| Dets, Seg, Ref, CL | UEMM-Air (Ours) | 2025 | ✓ | 6 | 120 | 13 | 1920×1080 | variable |
II Related Work
II-A UAV-based Environmental Perception
UAV-based environmental perception primarily involves two tasks: object detection and semantic segmentation.
II-A1 Object Detection
@@
Due to the typically top-down perspective of UAVs and the relatively small scale of the objects, general object detection methods [43, 44, 45, 46, 47] tend to be not suitable for UAV-based Object Detection (UAV-OD) tasks. The mainstream methods for UAV-based object detection primarily employ coarse-to-fine strategies [48, 49, 50]. Initially, the detection process focuses on identifying larger objects, while concurrently pinpointing dense subregions containing small objects. These subregions are subsequently utilized as inputs for the model to refine detection results. For example, a CZDetector [51] employed a density-based cropping algorithm to identify regions with crowded objects and then increased the size of those regions to enhance the training dataset. Alternatively, [52] utilizes a Gaussian mixture model to supervise the detector in generating object clusters composed of focusing regions. To address limited computing resources, methods were proposed to balance accuracy and efficiency. Typical examples include CEASC [53] and SIFDAL [37]. The former adopted a plug-and-play detection method with enhanced sparse convolution and an adaptive mask scheme. The latter disentangled scale-invariant features to boost detection accuracy and mildly reduce test inference costs. Additionally, to adapt to the low computational power conditions of UAVs, some researchers [54, 55] employed compress techniques such as pruning and distillation.
II-A2 Semantic Segmentation
Similar to object detection, there are many semantic segmentation methods specifically designed for UAV scenarios. High-resolution representation learning plays a crucial role in UAV semantic segmentation due to the ultra-high resolution and varying object scale of UAV remote sensing imagery [56, 57, 58]. For example, HRNet [59] was designed by a high-resolution network that repeatedly exchanges semantic information across adjacent multi-resolution sub-networks. At the same time, many lightweight methods were proposed to meet the real-time application requirements of UAV platforms [60, 61, 62]. Literature [63] also proposed Semantics Guided Bottleneck Network (SGBNet) based on BiSeNet and the Channel Pooled Attention (CPA) mechanism to balance segmentation accuracy, model size, and inference speed on the Land Cover Dataset.
II-B UAV-based Environmental Perception Datasets
@@
Many UAV-based environmental perception datasets provide multi-class images and videos captured by UAVs. These datasets are significant for promoting the research and development of various computer vision tasks, including object detection, object tracking, and scene understanding. We summarized several commonly used UAV-based datasets in Table I.
@@
Stanford-Drone [33] is a large-scale dataset containing overhead images and videos of multi-class objects moving and interacting at Stanford University. This dataset can be used for learning and evaluating multi-object tracking, activity understanding, and trajectory prediction.
@@
AU-AIR [26] includes extracted frames meta-data, bounding box annotations for traffic-related object categories, and multi-modal flight sensor data. The dataset is captured at low altitudes at the intersection.
@@
HIT-UAV [35] is the first publicly available high-altitude UAV-based infrared thermal dataset for object detection applications on UAVs. The dataset provides two types of bounding box annotations (oriented and standard) to tackle the challenge of overlapping object instances in aerial images.
RTDOD [36] is the first RGB-Thermal domain-incremental object detection dataset designed specifically for UAVs. The dataset covers a wide range of weather conditions and day-to-night transitions, highlighting the complexity of real-world scenarios.
State-Air [37] is an aerial dataset with multi-modal sensor data collected in real-world outdoor environments. The dataset concludes 2246 images of sunny days and 616 instances of snowy ones with 4 categories: person, car, bus, and van.
UAVid [38] is a semantic segmentation dataset designed for UAV semantic segmentation in complex urban scenes, featuring on both static and moving object recognition. It provides 300 high-resolution oblique-view UAV images, labeled with 8 classes, and gives a diverse representation of objects with rich scene context.
NWPU_YRCC [39] is the first public UAV image dataset containing 814 accurately annotated images for river ice segmentation. It covers typical images of river ice in different periods, with diverse appearances and captured from different flight heights and views.
RescueNet [40] propose a high-resolution post-disaster dataset for natural disaster damage assessment. It includes detailed classification and semantic segmentation annotations, enabling applications in building damage classification, road segmentation, and future disaster assessment.
AFID [41] is a publicly available dataset of aerial and fluvial images, featuring detailed semantic annotation from different camera perspectives. It focuses on utilizing semantic segmentation models to fulfill unmanned hydrologic data collection, environmental inspection, and disaster warning tasks.
SkyScenes [20] is a synthetic dataset of densely annotated aerial images captured from UAV perspectives, containing 33.6K images from different altitudes and pitch settings. It provides pixel-level semantic, instance, and depth annotations, and enables reproducing the same viewpoint under different conditions.
@@
III UEMM-Air
III-A Scene Construction and Flight Control Logic
Previous UAV-based datasets are limited in scene diversity, which tends to affect model generalization. Therefore, we aim to construct a dataset with richer scenes to improve the performance of the models. To be specific, we utilize Unreal Engine [31] with CityBLD [64] plugin. It can create cities of almost any size and style in a very short time to simulate scenarios in the real world. We build several scenes including cities, parks, highways, etc. We also collect a total of 13 categories and more than hundreds of vehicle models.
@@ III-A Scene Construction and Flight Control Logic
To collect data, we control the UAV to fly and take pictures in Unreal Engine, as represented in Fig. 3 Step 1. Specifically, we build an Unmanned Aerial Vehicle simulator using AirSim and Pygame. When flying to a satisfactory shooting point, we control the UAV to fly within a height range of 5 meters to 50 meters above the horizontal surface, taking a set of photos every 5 meters up. The camera rotates from 0 degrees to 90 degrees by 5 degrees for each step. To obtain aligned pictures of different modalities, our simulator temporarily stops running when taking photos.
Ultimately, we built 8 different maps and collected a total of pairs of data. Each map has its own unique style, incorporating various elements such as different lighting, scenes, and weather conditions. As shown in Fig. 4, we randomly select 1,000 images from each map and utilize CLIP to extract image features for T-SNE visualization analysis. The results indicate considerable feature differences among different maps, providing rich domain representations for model training.
III-B Acquisition Setup
@@
III-C Automatic Image Segmentation and Annotation
Most of the existing UAV-based datasets are manually annotated. Manual image annotation faces challenges in terms of accuracy and efficiency, especially when dealing with a large number of labels or low-resolution images.
To avoid manual annotation, the SynDrone [65] dataset employs an automatic image labeling algorithm. They derive the absolute coordinates of UAV and vehicles from Unreal Engine, then obtain the bounding boxes of the objects by analyzing their relative position. However, this strategy causes some incorrect annotations where objects are visibly blocked but their coordinates are still marked on the image, as illustrated in Fig. 5.
To avoid mislabeling visually overlapped objects, we utilize depth images where pixel value represents the distance from the object to the camera plane to perform a secondary annotation. Intuitively, depth values mildly change on each object and a depth value jump indicates multiple objects existences. Therefore, overlapped objects can be correctly identified through depth observation. We detect depth mutations within segmented bounding boxes to confirm object edges, adjusting labels accordingly.
III-D Cross-modality Generation
After collecting the 5 modalities mentioned in Section III-B, we also need to generate the sixth modality: text. In this paper, we categorize text into two types: global and local, corresponding to the tasks of image-text contrastive learning and referring image segmentation, respectively.
III-D1 Image-text Contrastive Learning
In addition to visual tasks like object detection and instance segmentation, we also hope our UEMM-Air can support vision language tasks to make UAV-based models perform zero-shot capability. In Section III-C, we already generate accurate annotations with object bounding boxes and fine-grained class names. However, such annotations cannot be directly utilized on the text encoder of CLIP. Therefore, we follow the B2C method from RemoteCLIP [66] to transfer the bboxes into a set of natural language captions. However, we notice that the captions generated by RemoteCLIP only include the quantity and position of the objects relative to the image, leading to insufficient semantic information in the captions. To alleviate this issue, we propose a new generation approach, as shown in Fig. 3. Specifically, we combine the scene and UAV parameter annotations provided by UEMM-Air to generate text descriptions that are more relevant to the UAV context. In addition, we provide more precise location information. Instead of being limited to the center and edges of the image, we utilize multiple sentences to comprehensively describe the distribution of the objects.
Ultimately, we generated 7 distinct captions for each image, resulting in a total of 840,000 descriptions. We present a visualization of the caption length distribution for our final data, as represented in Fig. 7. It can be observed that the caption length distribution shows a peak around 200, with most lengths concentrated between 200-400. Beyond 400, the frequency decreases, forming a long-tail pattern. In Fig. 8, we also provide visualizations of word clouds and the top 20 keywords of our UEMM-Air. The words exclude stop words like ”the”, ”is”, and others.
III-D2 Referring Image Segmentation
As a prominent visual-language task, referring image segmentation enables the delineation of specific objects in the visual field through natural language descriptions. This capability is equally essential in the domain of UAVs to facilitate cross-modal interactive tasks, thereby advancing embodied intelligence in UAVs. Therefore, similar to image-text contrastive learning, we propose a method for generating textual descriptions for the referring image segmentation task. By analyzing the spatial relationships of the objects within the obtained segmentation labels, we can automate the generation of the referring captions.
Unlike referring image segmentation tasks in general domains, the UAV field often requires the segmentation of multiple objects rather than just one. For instance, it may involve segmenting a series of cars along a street. Consequently, we primarily generate descriptions in the following 3 categories:
All objects of a specific class. e.g., All of the BoxTrucks in the image.
The spatial relationships of a particular object or certain objects relative to the image. e.g., Van in the bottom-left region of the image.
The spatial relationships of a specific object relative to other objects. e.g., The Van above the SUV in the image.
Finally, we generated 600,000 image-mask-text pairs. Additionally, we present the statistics on captions length and word clouds in Fig. 7 and Fig. 8, respectively. Unlike the data generated in image-text contrastive learning, the length of the captions for referential tasks is shorter, as it does not require a description of the global features of the image.
III-D3 Visualizations of Cross-modality Generation
To validate the accuracy of the generated captions, we randomly selected samples for visualization. As shown in Fig. 9, the captions for the image-text retrieval task accurately describe the number and spatial relationships of the objects, as well as the UAV’s height and angle information. The captions for the referring image segmentation specifically describe the location of a particular object and successfully generate the mask for the object.
IV Benchmark and Experiments
IV-A Experimental Setup
IV-A1 Benchmark Models